OpenStudy (anonymous):

y=2x^2+12x+1 in vertex form

OpenStudy (immanuelv):

do you know vertex form?

OpenStudy (anonymous):

no it just says write

OpenStudy (immanuelv):

The vertex form of a quadratic is given by y = a(x – h)2 + k, where (h, k) is the vertex.

OpenStudy (immanuelv):

y = a(x – h)^2 + k, where (h, k) is the vertex.

OpenStudy (immanuelv):

y = a(x-h)^2 +k

OpenStudy (immanuelv):

does that help?